# Bayesian optimization of electron energy from laser wakefield accelerator

### v2

@@ Abstract
We use Bayesian optimization in combination with three-dimensional particle-in-cell simulations to determine the optimal laser and plasma parameters that, for a given laser pulse energy, maximize the cut-off energy of an electron beam produced by laser wakefield accelerators. We assume a Gaussian laser driver with matched spot size and amplitude and investigate both self-guiding in a uniform-density plasma and guiding in a preformed plasma channel with matched radius. To interpret the simulation results quantitatively, we derive novel analytical expressions for the maximum electron energy and the corresponding acceleration distance, accounting for the effects of laser diffraction and energy depletion. Based on the results obtained, we discuss the potential scalability of the optimal input (plasma density, pulse duration, amplitude, and spot size) and output (electron energy and acceleration length) parameters to laser systems of arbitrary energy.
@@ I Introduction
In this work, we aim to identify the optimal set of laser and plasma parameters that, for a given laser pulse energy, maximize the cut-off energy (i.e., the energy threshold beyond which there are very few or no particles) of an electron beam accelerated via LWFA. At this stage, we focus solely on optimizing the cut-off energy, without regard to other electron beam parameters. This focus is motivated by the needs of specific applications where achieving a cut-off energy above certain threshold is essential (e.g., muon production [11, 12, 13] and nuclear activation [14, 15]).
@@ II Propagation of intense laser pulse in plasma
where $a_{0}$ is the strength parameter of the laser pulse and $w_{0}$ is the laser waist (i.e., the radius at which the intensity drops to $1/\mathrm{e}^{2}$ of its peak value at the focus). The Rayleigh length, $x_{R}=k_{x}w_{0}^{2}/2$, is defined as the distance from the focus along the propagation direction to the point where the laser radius increases by a factor of $\sqrt{2}$ compared to the waist.
@@
are given. Here, ¯$\overline{\mathcal{E}}=\pi\overline{\mathcal{P}}/\omega_{0}$, $\overline{\mathcal{P}}=2m_{e}c^{3}/r_{e}$, $r_{e}=e^{2}/4\pi\epsilon_{0}m_{e}c^{2}$ is the classical electron radius, and $\epsilon_{0}$ is the vacuum permittivity. The laser power is then
@@
$\approx 4.84\times 10^{-26}\,\mathrm{\frac{m\cdot s}{J}}\,\frac{\lambda_{0}^{2}\mathcal{E}_{0}n_{e}}{\tau_{0}}.$
@@
$\approx 7.29\times 10^{-9}\,\left(\mathrm{\frac{m\cdot s}{J}}\right)^{1/3}\,\left(\frac{\lambda_{0}^{2}\mathcal{E}_{0}n_{e}}{\tau_{0}}\right)^{1/3}$
@@
$\approx 9.07\times 10^{2}\,\left(\mathrm{\frac{s}{J\cdot m^{2}}}\right)^{1/6}\,\left(\frac{\lambda_{0}^{2}\mathcal{E}_{0}}{n_{e}^{2}\tau_{0}}\right)^{1/6}.$
@@
$R_{\mathrm{ch}}\left[\upmu\mathrm{m}\right]\approx 0.17\left(w_{0}\left[\upmu\mathrm{m}\right]\right)^{2}\sqrt{n_{e}\left(0\right)\left[10^{19}\,\mathrm{cm}^{-3}\right]}.$ (14)
@@ III.1 Unguided pulse
×$\times\exp{\left[-\frac{k_{p}^{2}l_{0}^{2}}{32\ln{2}}-k_{p}x_{R}\left(1-\beta_{\mathrm{ph}}\right)\right]}.$
@@ III.2 Guided pulse
×$\times\frac{2\cos{\phi}-k_{p}l_{pd}\left(1-\beta_{\mathrm{ph}}\right)\sin{\phi}}{4+k_{p}^{2}l_{pd}^{2}\left(1-\beta_{\mathrm{ph}}\right)^{2}}.$
Taking into account the pump depletion length in the form $l_{pd}\approx l_{0}k_{0}^{2}/k_{p}^{2}$ [46], Eq. (23) reaches its maximum for $k_{p}l_{0}\approx 4.07$ (i.e., $\tau_{0}>\tau_{r}$) and $\phi\approx-0.79$. Assuming further that $\gamma_{\mathrm{ph}}\gg 1$ (i.e., the low-density plasma limit), the expression for the maximum electron energy in the case of a guided laser pulse simplifies to
@@ IV Bayesian optimization
Kernels (also referred to as covariance functions) are a key component that determines the shape of both the prior and posterior distributions of the GP. We use the Matérn kernel, which provides a good balance between smoothness and flexibility for modeling non-linear objective landscapes typical of LWFA optimization. The Matérn kernel includes a parameter $\nu$ that controls the smoothness of the resulting function. We chose $\nu=5/2$, meaning the learning function is (at least) twice differentiable.
A critical aspect of BO is balancing exploration (i.e., searching the broader parameter space) and exploitation (i.e., probing areas near the current known optimum). The UCB acquisition function includes a free parameter that allows us to control this balance. We set this parameter to be decaying, i.e., it initially prioritizes exploration to allow a thorough search of the parameter space, and gradually shifts towards exploitation as more points are evaluated. This strategy minimizes the risk of getting trapped in local extrema and increases the likelihood of finding the global optimum.
@@
$n_{e}\approx 2.07\times 10^{39}\,\mathrm{m^{-3}\cdot s^{-1}}\,\tau_{0}$ (26)
@@
$\tau_{0}\approx 4.17\times 10^{-2}\,\mathrm{m^{-3/2}\cdot s}\,n_{e}^{-1/2},$ (27)
respectively, intersect at $n_{e}\approx 1.95\times 10^{19}\,\mathrm{cm^{-3}}$ and $\tau_{0}\approx 9.44\,\mathrm{fs}$. At the intersection, the matching conditions (11) and (12) yield $a_{0}=2$ and $w_{0}\approx 3.4\,\upmu\mathrm{m}$, respectively. We restrict the parameter space for the BO algorithm to $\tau_{0}\in[3,12]\,\mathrm{fs}$ and $n_{e}\in[1,4]\times 10^{19}\,\mathrm{cm^{-3}}$. If these bounds prove inadequate, they can be dynamically adjusted during the optimization process.
While we did not implement an adaptive stopping criterion, we monitored the relative improvement in the objective function over the last several evaluations. We terminated the optimization when (i) a sufficiently dense set of evaluated points existed around the current optimum candidate (i.e., the surrounding space was well sampled in all directions), (ii) the acquisition function near the candidate maximum was low, indicating minimal uncertainty or potential improvement, and (iii) additional samples around the current maximum did not significantly affect its value or position.
Figure 1: (Color online). (a) Ratio of laser power to critical power for relativistic self-focusing, $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$, (b) ratio of pulse duration to resonant duration, $\tau_{0}/\tau_{r}$, (c) strength parameter, $a_{0}$, and (d) beam waist, $w_{0}$, given by Eqs. (9), (10), (11) and (12), respectively. All variables are displayed within the prescribed limits of electron density, $n_{e}$, and pulse duration, $\tau_{0}$, (or, equivalently, laser power, $\mathcal{P}_{0}$) considering laser central wavelength $1\,\mathrm{\upmu m}$ and energy $10\,\mathrm{mJ}$. The solid line in (a) and the dashed line in (b) show the contour curves of $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$ according to Eqs. (26) and (27), respectively.
Panels (a)–(d) of Fig. 1 show $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$, $\tau_{0}/\tau_{r}$, $a_{0}$, and $w_{0}$ according to Eqs. (9), (10), (11), and (12), respectively. All variables are displayed within the prescribed limits of $n_{e}$ and $\tau_{0}$, assuming $\lambda_{0}=1\,\upmu\mathrm{m}$ and $\mathcal{E}_{0}=10\,\mathrm{mJ}$. The contours of $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$, according to Eqs. (26) and (27), are shown in panels (a) and (b), respectively.
@@ V Particle-in-cell simulations
[…]
@@ VI Optimization results
In what follows, we define the cut-off energy of the accelerated electron beam, $\mathcal{E}_{e,\mathrm{max}}$, as the upper limit of its energy distribution with a spectral density of at least $10\,\mathrm{fC/MeV}$. The corresponding acceleration length, $l_{\mathrm{acc}}$, is defined as the distance from injection to the point where the electrons reach their maximum cut-off energy. We do not distinguish whether this length is limited by laser diffraction, energy depletion, electron dephasing, or other effects.
@@ VI.1 Self-guiding in uniform-density plasma
[…]
Figure 3: (Color online). Spatial and spectral characteristics of the output electron beam obtained from $A_{16}$ – the PIC simulation with parameters closest to the optimum found by BO in the case of self-guiding in a uniform-density plasma (parameters are listed in Tab. 1). (a) 3D visualization of electron density, $n_{e}/n_{\mathrm{cr}}$, laser intensity, $I/I_{0}$ (with $I_{0}=2\mathcal{P}_{0}/\pi w_{0}^{2}$), and the energy of accelerated electrons, $\mathcal{E}_{e}$, considering only electrons with $\mathcal{E}_{e}>10\,\mathrm{MeV}$. The electron density is sliced along the $x$$z$ plane to reveal the inner structure of the plasma wave. (b) (top) Energy spectrum of the electron beam, $\mathrm{d}|\mathcal{Q}|/\mathrm{d}\mathcal{E}_{e}$. (bottom left) Distribution of the electron charge with respect to energy and propagation angle, $\mathrm{d}^{2}|\mathcal{Q}|/\mathrm{d}\mathcal{E}_{e}\mathrm{d}\theta$. (bottom right) Angular spectrum of the electron beam, $\mathrm{d}|\mathcal{Q}|/\mathrm{d}\theta$.
@@
Panels (a) and (b) of Fig. 3 display the spatial and spectral characteristics, respectively, of the output electron beam obtained from $A_{16}$ – the PIC simulation with parameters closest to the optimum found by BO in a uniform-density plasma (parameters are listed in Tab. 1). As can be seen, the energy spectrum is relatively broad, as it includes electrons injected into multiple periods of the plasma wave, not only via nanoparticle-assisted injection but also through self-injection occurring later during the interaction. The total charge of electrons with energy above $1\,\mathrm{MeV}$ is $\approx 40\,\mathrm{pC}$, and their angular distribution shows a FWHM divergence of $\approx 37\,\mathrm{mrad}$. We note that these beam characteristics were not the subject of optimization.
[…]
Figure 5: (Color online). Spatial and spectral characteristics of the output electron beam obtained from $B_{17}$ – the PIC simulation with parameters closest to the optimum found by BO in the case of a preformed plasma channel with matched radius (parameters are listed in Tab. 2). (a) The 3D visualization of electron density, $n_{e}/n_{\mathrm{cr}}$, laser intensity, $I/I_{0}$, where $I_{0}=2\mathcal{P}_{0}/\pi w_{0}^{2}$, and the energy of accelerated electrons, $\mathcal{E}_{e}$, considering only electrons with $\mathcal{E}_{e}>10\,\mathrm{MeV}$. The electron density is sliced along the $x$$z$ plane to reveal the inner structure of the plasma wave. (b) (top) The energy spectrum of the electron beam, $\mathrm{d}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{E}_{e}$. (bottom left) The distribution of the electron charge with respect to the electron energy and propagation angle, $\mathrm{d}^{2}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{E}_{e}\mathrm{d}\theta$. (bottom right) The angular spectrum of the electron beam, $\mathrm{d}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{\theta}$.

### v3

@@ Abstract
We use Bayesian optimization in combination with three-dimensional particle-in-cell simulations to determine the optimal laser and plasma parameters that, for a given laser pulse energy, maximize the cut-off energy of an electron beam produced by laser wakefield accelerators. We assume a Gaussian laser driver with matched spot size and amplitude and investigate both self-guiding in a uniform-density plasma and guiding in a preformed plasma channel with matched radius. To interpret the simulation results quantitatively, we derive novel analytical expressions for the maximum electron energy and the corresponding acceleration distance, accounting for the effects of laser diffraction and energy depletion. Based on the results obtained, we discuss the potential scalability of the optimal input (plasma density, pulse duration, amplitude, spot size, and channel radius) and output (electron energy, electron charge, acceleration length, and acceleration efficiency) parameters to laser systems of arbitrary energy.
@@ I Introduction
In this work, we aim to identify the optimal set of laser and plasma parameters that, for a given laser pulse energy, maximize the cut-off energy (i.e., the energy threshold beyond which there are very few or no particles) of an electron beam accelerated via LWFA. At this stage, we focus solely on optimizing the cut-off energy, without regard to other electron beam parameters. This focus is motivated by the needs of specific applications where achieving a cut-off energy above certain threshold is essential (e.g., muon production [11, 12, 13, 14] and nuclear activation [15, 16]).
@@ II Propagation of intense laser pulse in plasma
where $a_{0}=eE_{0}/m_{e}\omega_{0}c$ is the strength parameter of the laser pulse (with $E_{0}$ being the electric field amplitude) and $w_{0}$ is the laser waist (i.e., the radius at which the intensity drops to $1/\mathrm{e}^{2}$ of its peak value at the focus). The Rayleigh length, $x_{R}=k_{x}w_{0}^{2}/2$, is defined as the distance from the focus along the propagation direction to the point where the laser radius increases by a factor of $\sqrt{2}$ compared to the waist.
@@
are given. Here,
$\overline{\mathcal{E}}=\frac{\pi\overline{\mathcal{P}}}{\omega_{0}}\quad\mathrm{and}\quad\overline{\mathcal{P}}=\frac{2m_{e}c^{3}}{r_{e}},$ (7)
$r_{e}=e^{2}/4\pi\epsilon_{0}m_{e}c^{2}$ is the classical electron radius, and $\epsilon_{0}$ is the vacuum permittivity. Note that $\overline{\mathcal{E}}\approx 29\ \mathrm{\upmu J}$ for $\lambda_{0}=1\ \mathrm{\upmu m}$ and $\overline{\mathcal{P}}\approx 17.4\ \mathrm{GW}$. The laser power is then
@@
[…]
@@
[…]
@@
[…]
@@
$R_{\mathrm{ch}}\approx 0.57w_{0}^{2}\sqrt{\pi r_{e}n_{e}\left(0\right)}.$ (15)
@@ III.1 Unguided pulse
$\times\exp{\left[-\frac{k_{p}^{2}l_{0}^{2}}{32\ln{2}}-k_{p}x_{R}\left(1-\beta_{\mathrm{ph}}\right)\right]}.$
@@ III.2 Guided pulse
$\times\frac{2\cos{\phi}-k_{p}l_{pd}\left(1-\beta_{\mathrm{ph}}\right)\sin{\phi}}{4+k_{p}^{2}l_{pd}^{2}\left(1-\beta_{\mathrm{ph}}\right)^{2}}.$
Taking into account the pump depletion length in the form $l_{pd}\approx l_{0}k_{0}^{2}/k_{p}^{2}$ [47], Eq. (24) reaches its maximum for $k_{p}l_{0}\approx 4.07$ (i.e., $\tau_{0}>\tau_{r}$) and $\phi\approx-0.79$. Assuming further that $\gamma_{\mathrm{ph}}\gg 1$, the expression for the maximum electron energy in the case of a guided laser pulse simplifies to
@@ IV Bayesian optimization
Kernels (also referred to as covariance functions) determine the shape of both the prior and posterior distributions of the GP. We use the Matérn kernel, which provides a good balance between smoothness and flexibility for modeling non-linear objective landscapes typical of LWFA optimization. The Matérn kernel includes a parameter $\nu$ that controls the smoothness of the resulting function. We chose $\nu=5/2$, meaning the learning function is (at least) twice differentiable.
A critical aspect of BO is balancing exploration (i.e., searching the broader parameter space) and exploitation (i.e., probing areas near the current known optimum). The UCB acquisition function includes a free parameter $\kappa$ that allows us to control this balance. We used a conservatively decaying $\kappa$ [50] to modestly shift the search toward exploitation in later iterations, while preserving some exploration capability given the limited evaluation budget.
@@
$n_{e}\approx 2.07\times 10^{39}\,\mathrm{m^{-3}\,s^{-1}}\,\tau_{0}$ (27)
@@
$\tau_{0}\approx 4.17\times 10^{-2}\,\mathrm{m^{-3/2}\,s}\,n_{e}^{-1/2},$ (28)
respectively, intersect at $n_{e}\approx 1.95\times 10^{19}\,\mathrm{cm^{-3}}$ and $\tau_{0}\approx 9.44\,\mathrm{fs}$. At the intersection, the matching conditions (12) and (13) yield $a_{0}=2$ and $w_{0}\approx 3.4\,\upmu\mathrm{m}$, respectively. We restrict the parameter space for the BO algorithm to $\tau_{0}\in[3,12]\,\mathrm{fs}$ and $n_{e}\in[1,4]\times 10^{19}\,\mathrm{cm^{-3}}$. If these bounds prove inadequate (i.e., when the current optimum lies on or very close to the search space boundary), they can be dynamically adjusted during the optimization process, allowing the optimizer to explore the surrounding region.
Figure 1: (Color online). (a) Ratio of laser power to critical power for relativistic self-focusing, $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$, (b) ratio of pulse duration to resonant duration, $\tau_{0}/\tau_{r}$, (c) strength parameter, $a_{0}$, and (d) beam waist, $w_{0}$, given by Eqs. (10)–(13), respectively. All variables are displayed within the prescribed limits of electron density, $n_{e}$, and pulse duration, $\tau_{0}$, (or, equivalently, laser power, $\mathcal{P}_{0}$) considering laser central wavelength $1\,\mathrm{\upmu m}$ and energy $10\,\mathrm{mJ}$. The solid line in (a) and the dashed line in (b) show the contour curves of $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$ according to Eqs. (27) and (28), respectively.
While we did not implement a specific adaptive stopping criterion (see, e.g., Refs. 51 and 52), we monitored the relative improvement in the objective function over the last several evaluations. The optimization was terminated when (i) a sufficiently dense set of evaluated points had been obtained around the current optimum candidate (i.e., the surrounding space was well sampled in all directions) and (ii) additional samples in this region did not significantly affect the estimated value or location of the optimum.
Panels (a)–(d) of Fig. 1 show $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$, $\tau_{0}/\tau_{r}$, $a_{0}$, and $w_{0}$ according to Eqs. (10)–(13), respectively. All variables are displayed within the prescribed limits of $n_{e}$ and $\tau_{0}$, assuming $\lambda_{0}=1\,\upmu\mathrm{m}$ and $\mathcal{E}_{0}=10\,\mathrm{mJ}$. The contours of $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$, according to Eqs. (27) and (28), are shown in panels (a) and (b), respectively.
@@ V Particle-in-cell simulations
We perform two sets of PIC simulations in 3D Cartesian geometry using the EPOCH code [53]. The laser and plasma parameters used in the simulations are listed in Tabs. 1 and 2 in the Appendix, corresponding to laser self-guiding in a uniform-density plasma and guiding in a preformed plasma channel with matched radius, respectively. The first four simulations in each case, $A_{1}$$A_{4}$ and $B_{1}$$B_{4}$, are selected based on an educated guess, while the subsequent simulations are determined iteratively by the BO algorithm.
@@ VI Optimization results
In what follows, we define the cut-off energy of the accelerated electron beam, $\mathcal{E}_{e,\mathrm{max}}$, as the upper limit of its energy distribution with a spectral density of at least $10\,\mathrm{fC/MeV}$. The corresponding acceleration length, $l_{\mathrm{acc}}$, is defined as the distance from injection to the point where the electrons reach their maximum cut-off energy. We do not distinguish whether this length is limited by laser diffraction, energy depletion, electron dephasing, or other effects. A single Bayesian optimization run was performed for each case (self-guiding in uniform-density plasma and guiding in preformed plasma channel with matched radius), and variability across independent runs was therefore not quantified.
@@ VI.1 Self-guiding in uniform-density plasma
Figure 2: (Color online). Self-guiding in a uniform-density plasma. (a) The GP model for the electron beam cut-off energy, $\mathcal{E}_{e,\mathrm{max}}$, and (b) the corresponding acceleration distance, $l_{\mathrm{acc}}$, based on the results of 18 PIC simulations, $A_{1}$$A_{18}$ (parameters are listed in Tab. 1). The solid and dashed lines in both panels correspond to $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$, respectively, as defined by Eqs. (27) and (28).
Figure 3: (Color online). Spatial and spectral characteristics of the output electron beam obtained from $A_{16}$ – the PIC simulation with parameters closest to the optimum found by BO in the case of self-guiding in a uniform-density plasma (parameters are listed in Tab. 1). (a) 3D visualization of electron density, $n_{e}/n_{\mathrm{cr}}$, laser intensity, $I/I_{0}$ (with $I_{0}=2\mathcal{P}_{0}/\pi w_{0}^{2}$), and the energy of accelerated electrons, $\mathcal{E}_{e}$, considering only electrons with $\mathcal{E}_{e}>10\,\mathrm{MeV}$. The electron density is sliced along the xxzz plane to reveal the inner structure of the plasma wave. (b) (top) Energy spectrum of the electron beam, $\mathrm{d}|\mathcal{Q}|/\mathrm{d}\mathcal{E}_{e}$. (bottom left) Distribution of the electron charge with respect to energy and propagation angle, $\mathrm{d}^{2}|\mathcal{Q}|/\mathrm{d}\mathcal{E}_{e}\mathrm{d}\theta$. (bottom right) Angular spectrum of the electron beam, $\mathrm{d}|\mathcal{Q}|/\mathrm{d}\theta$.
@@
Panels (a) and (b) of Fig. 3 display the spatial and spectral characteristics, respectively, of the output electron beam obtained from $A_{16}$ – the PIC simulation with parameters closest to the optimum found by BO in a uniform-density plasma (parameters are listed in Tab. 1). As can be seen, the energy spectrum is relatively broad, as it includes electrons injected into multiple periods of the plasma wave, not only via nanoparticle-assisted injection but also through self-injection occurring later during the interaction. The total charge of electrons with energy above $1\,\mathrm{MeV}$ is $\approx 40\,\mathrm{pC}$, their angular distribution shows a FWHM divergence of $\approx 37\,\mathrm{mrad}$, and the efficiency of energy transfer from the laser to the accelerated electron beam is $\approx 9.8\,\%$. We note that these beam characteristics were not the subject of optimization.
Figure 4: (Color online). Guiding in a preformed plasma channel. (a) The GP model for the electron beam cut-off energy, $\mathcal{E}_{e,\mathrm{max}}$, and (b) the corresponding acceleration distance, $l_{\mathrm{acc}}$, based on the results of 18 PIC simulations, $B_{1}$$B_{18}$ (parameters are listed in Tab. 2). The solid and dashed lines in both panels correspond to $\mathcal{P}_{0}=\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}=\tau_{r}$, respectively, as defined by Eqs. (27) and (28).
Figure 5: (Color online). Spatial and spectral characteristics of the output electron beam obtained from $B_{17}$ – the PIC simulation with parameters closest to the optimum found by BO in the case of a preformed plasma channel with matched radius (parameters are listed in Tab. 2). (a) The 3D visualization of electron density, $n_{e}/n_{\mathrm{cr}}$, laser intensity, $I/I_{0}$, where $I_{0}=2\mathcal{P}_{0}/\pi w_{0}^{2}$, and the energy of accelerated electrons, $\mathcal{E}_{e}$, considering only electrons with $\mathcal{E}_{e}>10\,\mathrm{MeV}$. The electron density is sliced along the xxzz plane to reveal the inner structure of the plasma wave. (b) (top) The energy spectrum of the electron beam, $\mathrm{d}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{E}_{e}$. (bottom left) The distribution of the electron charge with respect to the electron energy and propagation angle, $\mathrm{d}^{2}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{E}_{e}\mathrm{d}\theta$. (bottom right) The angular spectrum of the electron beam, $\mathrm{d}\left|\mathcal{Q}\right|/\mathrm{d}\mathcal{\theta}$.